What is the resistance in () of an 18.5 m long piece of 29 gauge copper wire having a 0.286 mm diameter? X2
Answer #1

given L: 18.5 m , 8= 0.286/2 = 0.143 X10 m Pl Resistivity of Copper) = 1.68x108 rm A = Ar2 . * * 10.143 X10-3)2 = 6.42 X10-8